# Moiré Superradiance in Cavity Quantum Electrodynamics with Quantum Atom Gas

### v2

@@ Abstract
As a novel platform for exploring exotic quantum phenomena, the moiré lattice has garnered significant interest in solid-state physics, photonics, and cold atom physics. While moiré lattices in two- and three-dimensional systems have been proposed for neutral cold atoms, the simpler one-dimensional moiré effect remains largely unexplored. We present a scheme demonstrating moiré effects in a one-dimensional cold atom-cavity coupling system, which resembles a generalized open Dicke model exhibiting superradiant phase transitions. We reveal a strong link between the phase transition critical point and the one-dimensional moiré parameter. Additionally, we derive the cavity field spectrum, connected to the dynamical structure factor, and showcase controlled atomic diffusion. This work provides a new route for testing one-dimensional moiré effects with cold atoms and open new possibility of moiré metrology.
@@
Two-dimensional (2D) moiré lattices engineered by stacking two 2D periodic layers with a relative twisting angle have emerged as an intriguing new experimental platform in solid-state physics and optics, in which many exotic phenomena have been unraveled, including unconventional superconductivity Bistritzer and MacDonald (2011); Cao et al. (2018); Tarnopolsky et al. (2019), quantum Hall effect Dean et al. (2013), non-Abelian gauge potentials San-Jose et al. (2012), and localization Wang et al. (2020). As a powerful playground for quantum simulation Bloch et al. (2008), efforts are pushing forward towards implementing moiré lattice via trapping neutral cold atoms in 2D and three-dimensional optical lattices González-Tudela and Cirac (2019); Salamon et al. (2020a, b); Luo and Zhang (2021); Meng et al. (2023); Wang et al. (2024), which essentially focus on the moiré physics of flat band. Noteworthy while one-dimensional (1D) superlattice have already been implemented in cold atoms to demonstrate localization Roati et al. (2008); Schreiber et al. (2015), the 1D moiré effects in which moiré parameter would play a vital role in the system quantum properties remain largely unexplored. Whilst in the counterpart electronic and optical systems, the research on 1D moiré effects are burgeoning Vu and Das Sarma (2021); Talukdar et al. (2022); Yu et al. (2023); Gonçalves et al. (2024); Li et al. (2025). Question naturally arises on whether one can observe moiré effects in a 1D setup with cold atoms.
@@
Here we propose a scheme to test such effects, specifically in a cold atom-cavity coupling system enabling superradiant phase transition. As sketched in Fig. 1, a cold atom gas composed of $N$ atoms of mass $m$ are trapped along the axis of a standing-wave optical resonator in the $x$-direction, and also illuminated by a laser incident from the cavity side. The driving laser is detuned far below the atomic transition, thus are scattered by the atoms into the cavity mode with a pump strength $\varepsilon_{p}$. The system can be mapped to the well-known Dicke model in quantum optics Baumann et al. (2010); Nagy et al. (2010); Klinder et al. (2015a). Upon driving the pump across a critical value, the interplay between cavity-mediated global interaction among atoms and cavity decay would bring the system into a steady state with diverging excitations, in which photons are collectively scattered into the cavity mode and the atoms self-organize themselves in the combined net potential of trapping and emerging intracavity standing-wave potential. A lot of studies on systems of this type have been reported, both in theory Nagy et al. (2011); Sandner et al. (2015); Maschler et al. (2007); Cosme et al. (2018); Zheng and Cooper (2018); Yin et al. (2020); Mivehvar (2024); Han et al. (2024) and experiment Landig et al. (2015); Klinder et al. (2015b); Landig et al. (2016); Landini et al. (2018); Vaidya et al. (2018).
@@
Figure 1: Schematic diagram for dissipative phase transition in a cavity-assisted moiré lattice. Atoms are trapped by a optical lattice potential $V_{l}\cos^{2}(k_{l}x)$ along the cavity axis. Atoms scatter the side pump laser $\varepsilon_{p}$ into the cavity mode $\cos(k_{c}x+\phi)$. The inset show atom self-ordering in the combined 1D moiré lattice. Cavity decay rate is $\kappa$.
@@
To build the moiré lattice, we consider an additional one-dimensional (1D) optical lattice is applied on the atoms, the optical lattice potential $V_{l}\cos^{2}(k_{l}x)$ is chosen incommensurate to the intracavity standing-wave mode $\cos(k_{c}x+\phi)$. Then ratio $r=k_{c}/k_{l}$ defines the moiré ratio of a combined 1D bichromatic moiré lattice Vu and Das Sarma (2021). Specifically we set $r$ as the ratio of two consecutive numbers of the Fibonacci sequence, i.e., $r=f_{n+1}/f_{n}$. Fibonacci sequence is defined by the recursion relation $f_{n+1}=f_{n-1}+f_{n}$ with $f_{0}=f_{1}=1$, in the $n\rightarrow\infty$ limit $r$ approaches the golden ratio $(\sqrt{5}+1)/2$. The 1D moiré lattice is quasiperiodic with a unit cell length $L=2\pi/k_{0}$, where $k_{0}$ denotes the elementary wave number with $k_{l(c)}=f_{n(n+1)}k_{0}$. Here we introduce the lowest common numerator of $r$ as the moiré parameter $M$ ($=f_{n+1}$), in the following we illustrate with three cases of $M=1,3,5$, correspond to $n=1,3,4$ respectively. For simplicity we assume the relative phase $\phi=0$.
@@
[…]
@@
where $c=(c_{0},c_{1},\dots,c_{n_{c}})^{T}$, the recoil frequency $\omega_{k}=\hbar k_{0}^{2}/2m$, $\mathds{1}$ is the identity matrix, the matrices $\mathcal{K}$, $\mathcal{M}_{f_{n}}$ are given in sm . Since in real experiment cavity mode is usually fixed, then in the following we scale the energy with $\omega_{r}=\hbar k_{c}^{2}/2m=M^{2}\omega_{k}$. That is to say, moiré parameter is manipulated by $k_{l}$. Hamiltonian (4) typically represents an extended Dicke model, in which the quantum light field is effectively coupled to multilevel transitions of many atoms. In addition to scattering within the cavity, the externally applied lattice allows for momentum transfers among atoms, giving rise to moiré effects in atom self-ordering and superradiant phase transition as will be illustrated below.
Mean-field solution and excitations.— In the $N\rightarrow\infty$ mean-field limit, the operators $a$, $c_{0(n)}$ split into
@@
where $\alpha$, $\psi_{0(n)}$ are scaled steady state expectations with $\delta a$, $\delta c_{0(n)}$ characterize the corresponding quantum fluctuations, $\mu$ is the chemical potential. By solving stationary equations sm , the cavity field amplitude $|\alpha|$ are plotted in Fig. 2 versus the scaled pumping strength $\eta=\sqrt{N}\varepsilon_{p}$.
Figure 2: (a) Scaled mean-field value $|\alpha|$ versus scaled pumping strength $\eta$ for different $M$. The black solid line is the case without the external optical lattice. The inset showcase steady state momentum distribution for $M=5$. (b) The eigenvalues imaginary parts of the lowest atom polariton excitation and incoherent cavity excitation versus $\eta$ for $M=5$. The vertical grey line indicate the analytical phase transition point sm . The steady states and excitations are calculated for $V_{l}=-1$, $NU_{0}=-80$, $\Delta_{c}=-100$ and $\kappa=20$.
@@
Without the external optical lattice potential, the present system exactly resemble the open Dicke model Nagy et al. (2011). In the thermodynamical limit the mean-field solutions would predict a critical pumping strength $\eta_{c}=\sqrt{-(\Delta_{c}^{2}+\kappa^{2})\omega_{r}/2\Delta_{c}}$ with effective detuning $\Delta_{c}=\delta_{c}-NU_{0}/2$, which separates the normal phase $\left\{\alpha=0,\psi_{0}=(1,0,0,\dots)^{T}\right\}$ from the superradiant phase with a finite $\alpha$. In Fig. 2(a) this case is indicated by a black solid line. With the increase of moiré parameter $M$, one can see that the corresponding critical pumping gradually decreases. Upon the onset of superradiant phase transition, in the steady state atoms disperse from the homogeneous state into the modulated states, as illustrated in the inset of Fig. 2(a) for $M=5$.
@@
The relation between the critical pumping strength and the moiré parameter $M$ can be understood via a study on the excitation properties sm : The moiré lattice provides an additional scattering channel of atoms in the $\cos(2k_{l}x)$ mode are scattered by the pump field into the cavity photon, resulting in atomic polariton excitations with eigenfrequency $w_{j=2f_{n}-f_{n+1}}$ which is much less than that in the usual scattering channel involving atoms in the homogeneous state. Increasing $M$ effectively equals enlargement of the lattice constant, which will decrease the energy gap and thus facilitating superradiant phase transition. In Fig. 2(b) we mark the analytic critical value for $M=5$ as the vertical line, which approximately coincide with the point at which the imaginary parts of the lowest atom polariton excitation frequency (black solid lines) reach zero. Incoherent cavity field excitation $\left\langle\delta a^{\dagger}\delta a\right\rangle$ (red-dashed line) also becomes divergent upon the onset of phase transition.
Cavity field spectrum.— The onset of phase transition can be indicated by the dynamic structure factor, which is the Fourier transform of the intracavity field correlator $S(\omega)\equiv\mathcal{F}\left(\left\langle a^{\dagger}\left(t_{s}+t\right)a\left(t_{s}\right)\right\rangle\right)$ ($t_{s}$ is the time for the system to reach steady state) and thus can be probed in experiment Landig et al. (2015); Agarwal (2006). We calculate $S(\omega)$ sm with the results presented in Fig. 3.
To map the dynamical structure factor, in the spectra plotted we have dropped the coherent part of $\omega=0$ Landig et al. (2015), which would become prominent upon the onset of superradiant phase transition as photons are collectively scattered into the cavity mode. The spectra are almost symmetric with respect to $\omega=0$ and their peaks come in pairs. Peaks of cavity spectrum actually reflect frequencies of atomic polariton mode excitation consisting of photonic and atomic parts sm . In the regime far below the threshold, the spectra peaks are located at $\omega\simeq\pm\omega_{r}$ and ±$\pm\omega_{k}=\pm\hbar(2k_{l}-k_{c})^{2}/2m$, reflecting that the cavity photons are scattered from the atoms in the homogeneous mode and those in the mode of $\cos(2k_{l}x)$, respectively. With the increase of pumping strength, the peaks in both pairs gradually move toward each other, the ±$\pm\omega_{k}$ peaks even merge at the critical point. This behavior is because the photonic components are becoming larger in the atomic polariton modes sm , and upon the phase transition, the lowest atomic polariton excitation frequency approaches zero, as also indicated in Fig. 2(b). Beyond the critical point, more pairs of peaks appear with the increase of the moiré parameter $M$, and the peaks of a pair tend to repel from each other when the pumping is continuously enhanced. The phenomena root in the fact that more atomic polariton modes are excited and their energies increase as well. The intracavity field spectra (dynamical structure factor) predicted here can serve as evidence of moiré effects on the superradiant phase transition.
Truncated Wigner approximation (TWA) simulation.— In order to verify whether the phenomena predicted above in the thermodynamical limit can really take place in a finite quantum system, here we simulate the dynamics in the open system depicted by the master equation (1). Due to the in principle unconstrained Hilbert space dimension of the cavity field in a generalized Dicke model and its global coupling to all the atomic modes, a full quantum simulation hiring Monte-Carlo wavefunction (quantum jump) algorithm is usually limited to small system size Sandner et al. (2015); Halati et al. (2020). We adopt TWA to study the dynamics sm . The TWA method Sinatra et al. (2002); Blakie et al. (2008); Polkovnikov (2010); Schachenmayer et al. (2015) have found qualitative agreements with experimental results in a series of atom-cavity setup of time crystal experiments Keßler et al. (2021); Kongkhambut et al. (2021); Skulte et al. (2023).
@@
As shown in Fig. 4(b), in every run of TWA simulation we propagate an initial state of $N=1000$ atoms in the homogeneous mode $\psi_{0}$ while the other atomic modes and the cavity field are left empty, by ramping the cavity pumping up to a desired value $\eta$ and holding it there until a steady state is reached. Compared with mean-field results, the TWA calculation indicate that onset of superradiant phase transition will apparently takes place at a smaller critical pumping $\eta$, as illustrated in Fig. 4(a). This suggests that quantum fluctuations will lower the critical value for the phase transition to take place. Apart from that, the dependence of superradiant phase transition on the moiré parameter is verified. In the right panel of Fig. 4(a), for $\eta=0$ steady state population in homogeneous mode $\psi_{0}$ becomes smaller with increasing $M$. This behavior is because the external optical lattice scatter the $\psi_{0}$ atoms into $\psi_{2k_{l}}$, whose scaled kinetic energy $4k_{l}^{2}/k_{c}^{2}=4/r^{2}$ decreases when $M$ becomes larger. The larger seed population of $\psi_{2k_{l}}$ in turn triggers dynamical instability of the normal phase at a smaller pumping strength $\eta$ sm .
Figure 4: TWA simulation results with $N=1000$ obtained from sample of $500$ trajectory runs. (a) Steady state population in cavity mode (left panel) and homogeneous atomic mode $\psi_{0}$ (right panel) versus pumping strength $\eta$. The black solid line is the case without the external optical lattice. (b) Cavity field population dynamics at $M=3$ along with $\eta$ ramps to the value of $0.6\eta_{c}$. (c) Wigner distributions of cavity light field at the instant marked by the black dashed line in (b). The parameters are the same as those in Fig. 2, noticeably that TWA however predicts the superradiant phase transition to occur at lower $\eta$.
@@
Although the presented TWA results already showcase the properties of phase transition, it is argued that finite system size will smoothen the abrupt change upon phase transitions in the thermodynamical limit Sandner et al. (2015), rendering that neither $|\alpha|^{2}$ nor $|\psi_{0}|^{2}$ a good phase transition indicator. Phase transitions are accompanied with spontaneous symmetry breaking. In the absence of the external optical lattice, with the occurrence of superradiance and the formation of intracavity optical lattice, atoms can spontaneously choose even or odd lattice sites to reside, which entangle with cavity field of opposite amplitudes and result in a macroscopic superposition state of $|\text{even}\rangle\otimes|\alpha\rangle+|\text{odd}\rangle\otimes|-\alpha\rangle$ Domokos and Ritsch (2002); Maschler et al. (2007). Similar processes take place for the present moiré lattice. For the dynamics of above threshold pumping in the $M=3$ case studied in Fig. 4(b), we project the cavity field sample of $500$ trajectories at three different times onto Fig. 4(c), which unravel its Wigner distribution. The cavity field evolves from vacuum noise (left panel), then being stretched and splitted (middle panel), and finally forms a Schrödinger cat state with opposite amplitudes (right panel). The associated atom steady states would be that atoms located at a few different sites of the combined moiré lattice. Due to the inherent symmetry of steady state, $\langle a\rangle$ or equivalently $\langle c^{\dagger}\mathcal{M}_{f_{n+1}}c\rangle$ would give the value 00. To break the symmetry, one can project the system state with respect to one cavity state maximizing the Wigner distribution and resolve the corresponding order parameters Sandner et al. (2015); Ostermann et al. (2020).
Atomic diffusion.— Considering the long time for the system to reach steady state versus limited lifetime of cold atoms, on the atomic part it would be more practicable to observe moiré effects through their diffusion. The dependence of superradiant phase transition on the moiré parameter provides an extra control knob on atom diffusion. This can be implemented by loading atoms into the cavity and observing their transport along the moiré lattice. We study the atom diffusion by taking the initial atomic wavefunction to be a Gaussian wavepacket with width $w$, while the cavity is in vacuum. Then the pump is tuned on and the wavepacket width is estimated with $w=\sqrt{\left\langle\left(\Delta x\right)^{2}\right\rangle}$ sm .
@@
Figure 5: (a) Time scaling $\nu$ of atomic diffusion versus pumping strength $\eta$ for $M=3$ (green-dashed line) and $M=5$ (red-dashed line). The horizontal black-dashed line of $\nu=1/2$ indicates the critical scaling. (b) Atomic diffusion dynamics simulated for M31 ($M=3,\eta=0.46\eta_{c}$), M32 ($M=3,\eta=0.5\eta_{c}$), M51 ($M=5,\eta=0.4\eta_{c}$) and M52 ($M=5,\eta=0.46\eta_{c}$), which were also indicated in (a). The black line indicates the normal diffusion of $\nu=1/2$ that separates superdiffusion ($\nu>1/2$) and subdiffusion ($\nu<1/2$) regions. The parameters are the same as before.

### v3

@@ Abstract
As a novel platform for exploring exotic quantum phenomena, the moiré lattice has garnered significant interest in solid-state physics, photonics, and cold atom physics. While moiré lattices in two- and three-dimensional systems have been proposed for neutral cold atoms, the simpler one-dimensional moiré effect remains largely unexplored. We present a scheme demonstrating moiré effects in a one-dimensional cold atom-cavity coupling system, which resembles a generalized open Dicke model exhibiting superradiant phase transitions. We reveal a strong link between the phase transition critical point and the one-dimensional moiré parameter. Evidences of the one-dimensional moiré effect are explicitly explored, including cavity field spectrum, phase transition dynamics, and anomalous atomic diffusion. This work provides a new route for testing one-dimensional moiré effects with cold atoms and open new possibility of moiré metrology.
I introduction
@@
Two-dimensional (2D) moiré lattices engineered by stacking two 2D periodic layers with a relative twisting angle have emerged as an intriguing new experimental platform in solid-state physics and optics, in which many exotic phenomena have been unraveled, including unconventional superconductivity Bistritzer and MacDonald (2011); Cao et al. (2018); Tarnopolsky et al. (2019), quantum Hall effect Dean et al. (2013), non-Abelian gauge potentials San-Jose et al. (2012), and localization Wang et al. (2020). As a powerful playground for quantum simulation Bloch et al. (2008), efforts are pushing forward towards implementing moiré lattice via trapping neutral cold atoms in 2D and three-dimensional optical lattices González-Tudela and Cirac (2019); Salamon et al. (2020a, b); Luo and Zhang (2021); Meng et al. (2023); Wang et al. (2024), which essentially focus on the moiré physics of flat band. Noteworthy while one-dimensional (1D) superlattice have already been implemented in cold atoms to demonstrate localization Roati et al. (2008); Schreiber et al. (2015), the 1D moiré effects in which moiré parameter would play a vital role in the system quantum properties remain largely unexplored. Whilst in the counterpart electronic and optical systems, the research on 1D moiré effects are burgeoning Vu and Das Sarma (2021); Talukdar et al. (2022); Yu et al. (2023); Gonçalves et al. (2024); Li et al. (2025); Wang et al. (2025). Question naturally arises on whether one can observe moiré effects in a 1D setup with cold atoms.
@@
Here we propose a scheme to test such effects, specifically in a cold atom-cavity coupling system enabling superradiant phase transition. As sketched in Fig. 1, a cold atom gas composed of $N$ atoms of mass $m$ are trapped along the axis of a standing-wave optical resonator in the $x$-direction, and also illuminated by a laser incident from the cavity side. The driving laser is detuned far below the atomic transition, thus are scattered by the atoms into the cavity mode with a pump strength $\varepsilon_{p}$. The system can be mapped to the well-known Dicke model in quantum optics Baumann et al. (2010); Nagy et al. (2010); Klinder et al. (2015a). Upon driving the pump across a critical value, the interplay between cavity-mediated global interaction among atoms and cavity decay would bring the system into a steady state with diverging excitations, in which photons are collectively scattered into the cavity mode and the atoms self-organize themselves in the combined net potential of trapping and emerging intracavity standing-wave potential. A lot of studies on systems of this type have been reported, both in theory Nagy et al. (2011); Sandner et al. (2015); Maschler et al. (2007); Cosme et al. (2018); Zheng and Cooper (2018); Yin et al. (2020); Mivehvar (2024); Han et al. (2024); Habibian et al. (2013) and experiment Landig et al. (2015); Klinder et al. (2015b); Landig et al. (2016); Landini et al. (2018); Vaidya et al. (2018).
@@
Figure 1: Schematic diagram for dissipative phase transition in a cavity-assisted moiré lattice. Atoms are trapped by a optical lattice potential $V_{l}\cos^{2}(k_{l}x)$ along the cavity axis. Atoms scatter the side pump laser $\varepsilon_{p}$ into the cavity mode $\cos(k_{c}x+\phi)$. The inset shows that atoms are scattered into a few different modes in the combined 1D moiré lattice. Cavity decay rate is $\kappa$.
The paper is organized as follows: In Section II we introduce the model implementing moiré superradiance, which can be mapped onto an extended open Dicke model. The mean-field solution and excitation properties are studied in Section III, with which we unravel the 1D moiré effect and the physics beneath. In order to verify the moiré effect in a finite quantum system, we present numerical simulation using truncated Wigner approximation (TWA) in Section IV. Section V is devoted to the calculation of cavity field spectrum, which can be observed in experiment and serve as an moiré effect indicator. In addition to that, the moiré effect can be manifested via atomic wavepacket diffusion in the moiré potential, the anomalous atomic diffusion are explored in Section VI. Finally we summarize and outlook in Section VII.
II moiré superradiance and extended Dicke model
@@
To build the moiré lattice, we consider an additional one-dimensional (1D) optical lattice is applied on the atoms, the optical lattice potential $V_{l}\cos^{2}(k_{l}x)$ is chosen different from the intracavity standing-wave mode $\cos(k_{c}x+\phi)$. Then ratio $r=k_{c}/k_{l}$ defines the moiré ratio of a combined 1D bichromatic moiré lattice Vu and Das Sarma (2021). Specifically we set $r$ as the ratio of two consecutive numbers of the Fibonacci sequence, i.e., $r=f_{n+1}/f_{n}$. Fibonacci sequence is defined by the recursion relation $f_{n+1}=f_{n-1}+f_{n}$ with $f_{0}=0$ and $f_{1}=1$, in the $n\rightarrow\infty$ limit $r$ approaches the golden ratio $(\sqrt{5}+1)/2$. The 1D moiré lattice is quasiperiodic with a unit cell length $L=2\pi/k_{0}$, where $k_{0}$ denotes the elementary wave number with $k_{l(c)}=f_{n(n+1)}k_{0}$. Here we introduce the lowest possible integer numerator of $r$ as the moiré parameter $M$ ($f_{n+1}$), in the following we illustrate with three cases of $M=1,3,5$, correspond to $n=1,3,4$ respectively. For simplicity we assume the relative phase $\phi=0$.
@@
$H_{A}/\hbar$ $=\int dx\Psi^{\dagger}(x)\left[-\frac{\hbar}{2m}\partial_{x}^{2}+V_{l}\cos^{2}(k_{l}x)\right]\Psi(x),$
@@
where $c=(c_{0},c_{1},\dots,c_{n_{c}})^{T}$, the recoil frequency $\omega_{k}=\hbar k_{0}^{2}/2m$, $\mathds{1}$ is the identity matrix, the matrices $\mathcal{K}$, $\mathcal{M}_{f_{n}}$ are given in Appendix A. Since in real experiment cavity mode is usually fixed, then in the following we scale the energy with $\omega_{r}=\hbar k_{c}^{2}/2m=M^{2}\omega_{k}$. That is to say, moiré parameter is manipulated by $k_{l}$. Hamiltonian (4) typically represents an extended Dicke model, in which the quantum light field is effectively coupled to multilevel transitions. In addition to scattering within the cavity, the externally applied lattice allows for momentum transfers among atoms, giving rise to moiré effects in atom self-ordering and superradiant phase transition as will be illustrated below.
III mean field solution and excitations
In the $N\rightarrow\infty$ mean-field limit, the operators $a$, $c_{0(n)}$ split into
@@
where $\alpha$, $\psi_{0(n)}$ are scaled steady state expectations with $\delta a$, $\delta c_{0(n)}$ characterize the corresponding quantum fluctuations, $\mu$ is the chemical potential. By solving stationary equations containing $\alpha$ and $\psi_{0(n)}$ (see Appendix B) , the cavity field amplitude $|\alpha|$ are plotted in Fig. 2(a) versus the scaled pumping strength $\eta=\sqrt{N}\varepsilon_{p}$.
@@
Without the external optical lattice potential, the present system exactly resemble the open Dicke model Nagy et al. (2011). In the thermodynamical limit the mean-field solutions would predict a critical pumping strength $\eta_{c}=\sqrt{-(\Delta_{c}^{2}+\kappa^{2})\omega_{r}/2\Delta_{c}}$ with effective detuning $\Delta_{c}=\delta_{c}-NU_{0}/2$, which separates the normal phase $\left\{\alpha=0,\psi_{0}=(1,0,0,\dots)^{T}\right\}$ from the superradiant phase with a finite $\alpha$. In Fig. 2(a) this case is indicated by a black solid line. With the increase of moiré parameter $M$, one can see that the corresponding critical pumping gradually decreases. Upon the onset of superradiant phase transition, in the steady state atoms disperse from the homogeneous state into the modulated states in which atoms occupy modes of discrete momentum $p$, as illustrated in the inset of Fig. 2(a) for $M=5$.
The relation between the critical pumping strength and the moiré parameter $M$ can be understood via a study on the excitation properties presented in Appendix C , in which we derive the atom polariton excitation frequencies $w_{j}^{s}$, with whose lowest imaginary value approach 0 we determine the phase transition critical point, as shown in Fig. 2(b). In addition to that atoms in the homogeneous state scatter the pump field into the cavity mode and give rise to atom polariton excitation $w_{j=f_{n+1}}^{s}$, the moiré lattice provides an additional scattering channel hiring atoms in the $\cos(2k_{l}x)$ mode, resulting in atomic polariton excitations with eigenfrequency $w_{j=2f_{n}-f_{n+1}}^{s}$ whose absolute imaginary value is much smaller. Increasing $M$ effectively decrease the energy gap and thus facilitating superradiant phase transition. In Fig. 2(b) we mark the analytic critical value for $M=5$ as the vertical line, which approximately coincide with the point at which the imaginary parts of the lowest atom polariton excitation frequency ($j=2f_{n}-f_{n+1}$) reach zero. Incoherent cavity field excitation $\left\langle\delta a^{\dagger}\delta a\right\rangle$ (red-dashed line) also becomes divergent upon the onset of phase transition.
Figure 2: (a) Scaled mean-field value $|\alpha|$ versus scaled pumping strength $\eta/\eta_{c}$ for different $M$. The black solid line is the case without the external optical lattice. The inset present steady state momentum distribution for $M=5$. (b) The eigenvalues imaginary parts of atom polariton excitations versus $\eta$ are shown for two major scattering channels at $M=5$, in which the absolute values of the pair $j=2f_{n}-f_{n+1}$ is much smaller than those of $j=f_{n+1}$, thus determining the phase transition point. The pair $j=f_{n+1}$ would converge to 0 at around $\eta=\eta_{c}$ without the occurrence of phase transition, as indicated by the black-dashed lines. The vertical grey line indicate the analytical phase transition point. Red-dashed line showcase incoherent cavity excitation. The steady states and excitations are calculated for $V_{l}=-1$, $NU_{0}=-80$, $\Delta_{c}=-100$ and $\kappa=20$.
IV twa simulation
In order to verify whether the phenomena predicted above in the thermodynamical limit can really take place in a finite quantum system, here we simulate the dynamics in the open system depicted by the master equation (1). Due to the in principle unconstrained Hilbert space dimension of the cavity field in a generalized Dicke model and its global coupling to all the atomic modes, a full quantum simulation hiring Monte-Carlo wavefunction (quantum jump) algorithm is usually limited to small system size Sandner et al. (2015); Halati et al. (2020). We adopt TWA to study the dynamics with the details given in the supplemental material sm . The TWA method Sinatra et al. (2002); Blakie et al. (2008); Polkovnikov (2010); Schachenmayer et al. (2015) have found qualitative agreements with experimental results in a series of atom-cavity setup of time crystal experiments Keßler et al. (2021); Kongkhambut et al. (2021); Skulte et al. (2023).
Figure 3: TWA simulation results with $N=1000$ obtained from sample of 500500 trajectory runs. (a) Steady state population in cavity mode (left panel) and homogeneous atomic mode $\psi_{0}$ (right panel) versus pumping strength $\eta$. The black solid line is the case without the external optical lattice. (b) Cavity field population dynamics at $M=3$ along with $\eta$ ramps to the value of $0.6\eta_{c}$. (c) Wigner distributions of cavity light field at the instant marked by the vertical grey dashed line in (b). The parameters are the same as those in Fig. 2, noticeably that TWA however predicts the superradiant phase transition to occur at lower $\eta$.
@@
As shown in Fig. 3(b), in every run of TWA simulation we propagate an initial state of $N=1000$ atoms in the homogeneous mode $\psi_{0}$ while the other atomic modes and the cavity field are left empty, by ramping the cavity pumping up to a desired value $\eta$ and holding it there until a steady state is reached. Compared with mean-field results, the TWA calculation indicate that onset of superradiant phase transition will apparently takes place at a smaller critical pumping $\eta$, as illustrated in Fig. 3(a). This suggests that quantum fluctuations will lower the critical value for the phase transition to take place. Apart from that, the dependence of superradiant phase transition on the moiré parameter is verified. In the right panel of Fig. 3(a), for $\eta=0$ steady state population in homogeneous mode $\psi_{0}$ becomes smaller with increasing $M$. This behavior is because the external optical lattice scatter the $\psi_{0}$ atoms into $\psi_{2k_{l}}$, whose scaled kinetic energy $4k_{l}^{2}/k_{c}^{2}=4/r^{2}$ decreases when $M$ becomes larger. The larger seed population of $\psi_{2k_{l}}$ in turn triggers dynamical instability of the normal phase at a smaller pumping strength $\eta$.
@@
Although the presented TWA results already showcase the properties of phase transition, it is argued that finite system size will smoothen the abrupt change upon phase transitions in the thermodynamical limit Sandner et al. (2015), rendering that neither $|\alpha|^{2}$ nor $|\psi_{0}|^{2}$ a good phase transition indicator. Phase transitions are accompanied with spontaneous symmetry breaking. In the absence of the external optical lattice (Dicke model), with the occurrence of superradiance and the formation of intracavity optical lattice, atoms can spontaneously choose even or odd lattice sites to reside, which entangle with cavity field of opposite amplitudes and result in a macroscopic superposition state of $|\text{even}\rangle\otimes|\alpha\rangle+|\text{odd}\rangle\otimes|-\alpha\rangle$ Domokos and Ritsch (2002); Maschler et al. (2007). Similar processes take place for the present moiré lattice. For the dynamics of above threshold pumping in the $M=3$ case studied in Fig. 3(b), we project the cavity field sample of $500$ trajectories at three different times onto Fig. 3(c), which unravel its Wigner distribution. The cavity field evolves from vacuum noise (left panel), then being stretched and splitted (middle panel), and finally forms a Schrödinger cat state with opposite amplitudes (right panel). The associated atom steady states would be that atoms located at a few different sites of the combined moiré lattice. Due to the inherent symmetry of steady state, $\langle a\rangle$ or equivalently $\langle c^{\dagger}\mathcal{M}_{f_{n+1}}c\rangle$ would give the value 0. To break the symmetry, one can project the system state with respect to one cavity state maximizing the Wigner distribution and resolve the corresponding order parameters Sandner et al. (2015); Ostermann et al. (2020).
V cavity field spectrum
The onset of phase transition can be indicated by the dynamic structure factor, which is the Fourier transform of the intracavity field correlator $S(\omega)\equiv\mathcal{F}\left(\left\langle a^{\dagger}\left(t_{s}+t\right)a\left(t_{s}\right)\right\rangle\right)$ ($t_{s}$ is the time for the system to reach steady state) and thus can be probed in experiment Landig et al. (2015); Agarwal (2006). We calculate $S(\omega)$ in Appendix D with the results presented in Fig. 4.
To map the dynamical structure factor, in the spectra plotted we have dropped the coherent part of $\omega=0$ Landig et al. (2015), which would become prominent upon the onset of superradiant phase transition as photons are collectively scattered into the cavity mode. The spectra are almost symmetric with respect to $\omega=0$ and their peaks come in pairs. Peaks of cavity spectrum actually reflect frequencies of atomic polariton mode excitation consisting of photonic and atomic parts sm . In the regime far below the threshold, the spectra peaks are located at $\omega\simeq\pm\omega_{r}$ and $\pm\omega_{k}=\pm\hbar(2k_{l}-k_{c})^{2}/2m$, reflecting that the cavity photons are scattered from the atoms in the homogeneous mode and those in the mode of $\cos(2k_{l}x)$, respectively. With the increase of pumping strength, the peaks in both pairs gradually move toward each other, the $\pm\omega_{k}$ peaks even merge at the critical point. This behavior is due to that the photonic components are becoming larger in the atomic polariton modes. Upon the phase transition, the lowest atomic polariton excitation frequency approaches zero. These behaviors are consistent with those of atomic polariton excitations presented in Fig. 2(b). Beyond the critical point, more pairs of peaks appear with the increase of the moiré parameter $M$, and the peaks of a pair tend to repel from each other when the pumping is continuously enhanced. The phenomena root in the fact that more atomic polariton modes are excited and their energies increase as well. The intracavity field spectra predicted here can serve as evidence of moiré effects on the superradiant phase transition.
@@
Figure 5: (a) Time scaling $\nu$ of atomic diffusion versus pumping strength $\eta$ for $M=3$ (green-dashed line) and $M=5$ (red-dashed line). The horizontal black-dashed line of $\nu=1/2$ indicates the critical scaling. (b) Atomic diffusion dynamics simulated for M31 ($M=3,\eta=0.46\eta_{c}$), M32 ($M=3,\eta=0.5\eta_{c}$), M51 ($M=5,\eta=0.4\eta_{c}$) and M52 ($M=5,\eta=0.46\eta_{c}$), which were also indicated in (a). The black line indicates the normal diffusion of $\nu=1/2$ that separates superdiffusion ($\nu>1/2$) and subdiffusion ($\nu<1/2$) regions. Atoms are initially prepared in a Gaussian wavepacket with width $w=1/k_{c}$. The other parameters are the same as before.
VI anomalous diffusion dynamics
Considering the long time for the system to reach steady state versus limited lifetime of cold atoms, on the atomic part it would be more practicable to observe moiré effects through their diffusion. The dependence of superradiant phase transition on the moiré parameter provides an extra control knob on atom diffusion. This can be implemented by loading atoms into the cavity and observing their transport along the moiré lattice. We study the atom diffusion by taking the initial atomic wavefunction to be a Gaussian wavepacket with width $w$, while the cavity is in vacuum. Then the pump is tuned on and the wavepacket width is estimated with $w=\sqrt{\left\langle\left(\Delta x\right)^{2}\right\rangle}$ (see Appendix F for details).